OpenStudy (anonymous):

Please Help!! x/5=16/10

OpenStudy (espex):

If solving for x, multiply both sides by 5.

OpenStudy (anonymous):

so I get 1 16 80 5 10 50= 8/25

OpenStudy (anonymous):

???

OpenStudy (anonymous):

solve for proportion

OpenStudy (espex):

\[\frac {x}{5}=\frac{16}{10} \rightarrow x=\frac{16}{10}\times5 \rightarrow x=\frac{16}{2} \rightarrow x=8\]

OpenStudy (anonymous):

Thank You...can I divide 2 into 16 to get 8

OpenStudy (espex):

Yes

OpenStudy (anonymous):

x/5=16/10 10x=80 x=8